OpenStudy (anonymous):

Rewrite 1000^log_base10(x) as an expression of x without logs.

OpenStudy (asnaseer):

firstly, do you know this rule:\[\log_ab=c\implies b=a^c\]

OpenStudy (asnaseer):

?

OpenStudy (anonymous):

Sorry to go dark. I had to go to work. Forgot to log out. Yes, I do know that rule. I get that logs are the inverse of exponentials, but somehow the notation always seems to confuse me, and I'm not quite sure what to do with the log in the exponent.

ganeshie8 (ganeshie8):

hint : \(10^ {log_{10} x} = x \)

OpenStudy (anonymous):

So, when the x comes down from the log, why isn't it 10^x?

ganeshie8 (ganeshie8):

u familiar wid inverse functinons ?

ganeshie8 (ganeshie8):

if u let, f(x) = log x g(x) = 10^x then, g is the inverse of f so, g(f(x)) = x

OpenStudy (anonymous):

Yes. It's just something about the notation that makes it hard for me to get a handle on. Log is the inverse of exponential.

ganeshie8 (ganeshie8):

u take log of 'something', then exponent of the result, u get ur 'something' back

OpenStudy (anonymous):

For some reason this isn't clicking, though.

ganeshie8 (ganeshie8):

see it in two steps, first step u take log next step u take exponent

OpenStudy (anonymous):

x^10?

ganeshie8 (ganeshie8):

almost... but not quite correct \(\huge 1000^{\log_{10}(x)} \)

ganeshie8 (ganeshie8):

\(\huge (10^3)^{\log_{10}(x)} \)

ganeshie8 (ganeshie8):

see if u can simplify now :)

OpenStudy (anonymous):

The x comes down, and the base is the exponent, no?

OpenStudy (anonymous):

And then, I'm still not sure what to do with the thousand.

ganeshie8 (ganeshie8):

why would x come down ?

OpenStudy (anonymous):

Well, if x is the log, that means it it is the power of the base, right?

OpenStudy (anonymous):

I feel like I'm making this more complicated than it needs to be.

OpenStudy (anonymous):

...which is a recurring theme in my mathematical endeavors.

OpenStudy (anonymous):

log10(100)=2

ganeshie8 (ganeshie8):

noo ur logic is correct, but we're taking log base 10 for x, but in the bottom we have 1000, so x wont come down !

ganeshie8 (ganeshie8):

for x to come down, we need to have 10 in the bottom, not 1000

OpenStudy (anonymous):

Not that bottom. I'm not being precise enough, I guess.

ganeshie8 (ganeshie8):

\(10^{\log_{10}SOMETHING } = SOMETHING\)

OpenStudy (anonymous):

I should have said...out? 10 to the something is x?

ganeshie8 (ganeshie8):

\(\huge (10^3)^{\log_{10}(x)} \) how it would become, based on ur thinking ?

ganeshie8 (ganeshie8):

ok.. hey im getting late , need to go... hope u would get these cleared up by someone else... good luck cya

OpenStudy (anonymous):

Um, thanks.

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

um can any of yall help me on my ? please

OpenStudy (anonymous):

I can't even seem to get mine figured out. :-/

OpenStudy (anonymous):

mine is about my ipod touch 2g

OpenStudy (anonymous):

Then I think you're doing it wrong. This is a thread about Calculus.

OpenStudy (anonymous):

no its techly the geek section on open study lmfao

OpenStudy (anonymous):

http://www.cliffsnotes.com/study_guide/Logarithmic-Functions.topicArticleId-10792,articleId-10791.html this will help with your issue then lol now please help with mine

OpenStudy (asnaseer):

@double.emms - going back to my original statement above:\[\log_ab=c\implies b=a^c\]now, if you substitute the expression for c (i.e. \(c=\log_ab\)) into:\[b=a^c\]you end up with:\[b=a^c=a^{\log_ab}\]Hence:\[a^{\log_ab}=b\tag{1}\]Next I assume you also know this law of indicies:\[(x^a)^b=x^{ab}=(x^b)^a\tag{2}\] Now take a close look at your question:\[1000^{log_{10}x}=(10^3)^{log_{10}x}\]Make use of the relations in (1) and (2) and see if you can solve from here.